An academic person wants to assess whether her remedial tutoring has been effective for her five students. She decides to conduct related samples t-test and records the following grades for students prior to and after receiving her help.

Tutoring
Before After
2.5 3.1
2.6 3.0
3.1 3.7
3.0 3.2
2.8 3.6

(a) Test whether or not her tutoring is effective at a 0.05 level of significance. State the value of the test statistic. (Round your answer to three decimal places.)
t =

State the decision to retain or reject the null hypothesis.

(b) Compute effect size using estimated Cohen's d. (Round your answer to two decimal places.)
d =
Answer #1

Ans:

Before After d
1 2.5 3.1 0.6
2 2.6 3 0.4
3 3.1 3.7 0.6
4 3 3.2 0.2
5 2.8 3.6 0.8
d-bar 0.52
sd 0.2280

a)

Test statistic:

t=(0.52-0)/(0.228/SQRT(5))

t=5.100

df=5-1=4

p-value=tdist(5.1,4,1)=0.0035

As,p-value<0.05,we reject the null hypothesis.

b)effect size

Cohen's d=0.52/0.228=2.28